I'm 46 years old and am trying to refresh my algebra.
How would I graph the function y=f(x)=x^2


Answer by rfer(16322)

make an input/output table
------------------------
y and f(x) are the same thing,
y is a funtion of x.
------------
y=2^2
y=4
(2,4)
---------------
y=3^2
y=9
(3,9)
------------
y=4^2
y=16
(4,16)
--------------------
your graph will be curved
use some negative values for x to develope the porabola